X Company wants to estimate total costs (overhead plus administrative) in March, when production is expected to be 1,400 units. It analyzed January costs to determine the parameters of the overhead and administrative cost functions. January production was 1,000 units and costs were as follows:

Total overhead costs $68,000
Total fixed overhead costs 37,400
Total administrative costs 35,000
Total variable administrative costs 8,750


What is X Company's estimate of total costs in March [round unit costs to two decimal places]?
Answer #1

Total overhead costs = $68,000

Total fixed overhead costs = $37,400

Total variable overhead = Total overhead costs - Total fixed overhead costs

= 68,000 - 37,400

= $30,600

Variable overhead cost per unit = Total variable overhead/ Number of units

= 30,600/1,000

= $30.6

Total administrative costs = $35,000
Total variable administrative costs = $8,750

Total fixed administrative cost = Total administrative costs -Total variable administrative costs

= 35,000-8,750

= $26,250

Total variable administrative cost per unit = 8,750/1,000
= $8.75

Calculation of Total Costs in March
Total variable cost (30.6 x 1,400) 42,840
Total fixed overhead cost 37,400
Total variable administrative costs (8.75 x 1,400) 12,250
Total fixed administrative cost 26,250
Total Cost $118,740

Kindly give a positive rating if you are satisfied with this solution and please ask if you have any query.

Thanks